if f(x)=3x+2 find f(6) and f(-2)
Answer by tjnw79(57)

3x+2 is the same as 3(6)+2= when x=6
and 3x+2 is the same as 3(-2)+2=
So the first problem 3(6)+2 is 18+2=20
and 3(-2)+2 is -6+2=-4